### v1

# Logic-in-Frames: Dynamic Keyframe Search via Visual Semantic-Logical Verification for Long Video Understanding

Abstract
Understanding long video content is a complex endeavor that often relies on densely sampled frame captions or end-to-end feature selectors, yet these techniques commonly overlook the logical relationships between textual queries and visual elements. In practice, computational constraints necessitate coarse frame subsampling, a challenge analogous to “finding a needle in a haystack.” To address this issue, we introduce a semantics-driven search framework that reformulates keyframe selection under the paradigm of Visual Semantic-Logical Search. Specifically, we systematically define four fundamental logical dependencies: 1) spatial co-occurrence, 2) temporal proximity, 3) attribute dependency, and 4) causal order. These relations dynamically update frame sampling distributions through an iterative refinement process, enabling context-aware identification of semantically critical frames tailored to specific query requirements. Our method establishes new SOTA performance on the manually annotated benchmark in key-frame selection metrics. Furthermore, when applied to downstream video question-answering tasks, the proposed approach demonstrates the best performance gains over existing methods on LongVideoBench and Video-MME, validating its effectiveness in bridging the logical gap between textual queries and visual-temporal reasoning. The code will be publicly available.
Figure 1: Examples of four types of visual semantic-logical relationships in video QA detected by our VSLS: Time (text, time, pen), Attribute (man, attribute, white shirt), Spatial (copilot, spatial, Egyptian Pyramids), and Causal (man, causal, basketball). Green boxes indicate correct answers, while red boxes show baseline errors.
1 Introduction
Figure 2: Visual Semantic-Logical Search (VSLS) Framework for Efficient Keyframe Selection in Long Video Understanding. The figure illustrates our VSLS framework for efficient keyframe search in long video understanding tasks. The pipeline begins with sparse sampling from the input video, followed by iterative frame selection based on object detection and semantic logical relation verification. Key components include: 1) Query decomposition via LLM/VLM to extract cue objects, target objects, and four types of logical relations (spatial, temporal, attribute, and causal); 2) Adaptive grid sampling that selects frames based on an evolving confidence distribution; 3) Object detection using YOLO-World to identify query-relevant entities; 4) Multi-faceted confidence estimation combining object and logic relation confidence scores through a spline function; selected high-confidence frames are fed to downstream tasks.
Vision-Language Models (VLMs) [40] have achieved remarkable progress in visual understanding [47, 29], particularly in video understanding [32, 44], which holds significant potential for modeling real-world scenarios. However, effectively modeling raw video data remains challenging due to the high-dimensional nature of spatio-temporal information and inherent model complexity constraints. These limitations severely restrict temporal dependency modeling, confining existing approaches to processing only a limited number of frames simultaneously - a critical bottleneck for comprehensive video understanding. This paper addresses a pivotal research question: How can we efficiently and accurately select keyframes that are semantically critical for answering video-based queries?
We hypothesize that deconstructing visual semantic and logical cues (e.g., target objects, logical relations between visual entities) from textual queries enables effective identification of task-relevant frames through heuristic sampling and search. Building on this insight, we propose Visual Semantic-Logical Search (VSLS), a novel framework that iteratively selects minimal yet maximally informative frames through target object confidence estimation and joint verification of visual semantic logic. Our method achieves temporal localization through iterative optimization of the sampling distribution. This approach requires only sparse sampling (1.4%) to identify critical frames while maintaining the temporal and spatial coherence of visual entities, significantly reducing computational complexity compared to conventional dense sampling strategies.
Compared with existing methods, VSLS offers three unique advantages. First, VSLS enables highly efficient frame retrieval without dense captioning (e.g., [4, 15, 31]) or video clustering (e.g., [33, 24]). By iteratively refining a sampling distribution, VSLS achieves precise temporal localization with only sampling 1.4% frames on average of LVHaystack. Second, VSLS comprehensively accounts for logical relations in the query, going beyond simple target detection (e.g., [39]). Specifically, it supports four types of relational logic—spatial co-occurrence, temporal dependencies, attribute dependencies, and causal relations—allowing it to accommodate a broader range of query types. Third, VSLS functions as a plug-and-play module independent of both the query analysis LLM and the QA-oriented VLM, enabling integration into existing pipelines without interdependency among components.
We examined VSLS on several public datasets, including LongVideoBench [38], a comprehensive benchmark for evaluating models on long-form video understanding tasks; Video-MME [11], a widely adopted multimodal evaluation benchmark for video understanding; and Haystack-LVBench [38], which features meticulously annotated frame-level labels for precise temporal analysis. Our extensive experiments demonstrate significant improvements in the semantic similarity between our retrieved keyframes and ground truth, temporal coverage, and accuracy on downstream long video QA tasks. More importantly, by sampling only 1.4% of video frames (Ego4D [12]) for detection and updating the sampling distribution, we achieved an 8.7% improvement in GPT-4o [14]’s long video QA accuracy. This result is attributed to our simple yet powerful observation: query-guided visual semantic logic retrieval can mitigate the gap between potential visual logic and the logic expressed in the query text. These findings suggest that constructing ternary logic triplets with visual elements (e.g., object1, logic type, object2) can enhance downstream reasoning capabilities when performing text-visual retrieval.
To our best knowledge, we are arguably the first to search for keyframes in long videos by detecting visual semantic logic, with potential extensions to other text-vision retrieval tasks. To summarize, we make the following contributions:
  • We define four prevalent types of semantic logic relations in video QA tasks, including temporal, causal, attribute, and spatial relations, which can be accurately detected.
  • We require sampling only 1.4% of frames on average to complete the keyframe search through (1) heuristic sampling and (2) updating the sampling distribution according to different video semantics and logical relation.
  • We validate several metrics, including retrieval efficiency, semantic similarity, and temporal coverage on multiple widely used open-source datasets, demonstrating significant improvements in downstream tasks.
2 Method
Although existing long-context VLM frameworks implement keyframe search for video QA tasks [17, 21, 28, 31, 33, 41], their computational efficiency and searching accuracy remain suboptimal. To address this needle-in-a-haystack challenge [30, 46], we propose a novel method VSLS that aligns the semantic relations between the text modality and video modality, enhancing the plausibility of logical reasoning and performance of downstream tasks.
2.1 Task Formulation
Given a video sequence $V=\{f_{t}\}_{t=1}^{N_{v}}$ with $N_{v}$ frames and query $Q$, the temporal search aims to determine the minimal keyframe subset $V^{K}=\{f_{m_{i}}\}_{i=1}^{m}\subseteq V$ that satisfies:
  • Conservation: The keyframe subset $V^{K}\subseteq V$ must satisfy the answer consistency condition: $\mathcal{A}(V^{K},Q)=\mathcal{A}(V,Q)$, where $\mathcal{A}(\cdot)$ denotes the video QA function.
  • Compactness: $V^{K}$ must be a minimal subset preserving completeness, meaning no frame can be removed without affecting answer consistency, also ensuring efficiency.
2.2 Visual Semantic Logic Extraction
Given a question $Q$ and uniform sampled frames $\overline{V}_{N}$ from a video $V$, our goal is to identify and organize the essential visual elements for answering $Q$. To this end, we categorize detected items into two groups:
  • Key Objects: The main participants or references in the scene that the question explicitly or implicitly focuses on (e.g., “person”, “microphone”).
  • Cue Objects: Secondary or contextual entities that help locate or disambiguate the Key Objects (e.g., “book”, “tiger painting”).
To reason about semantic and logical links among these objects, we define a set of relations $\mathcal{R}\subseteq\mathcal{O}\times\Delta\times\mathcal{O}$, where each relation $r\in\mathcal{R}$ is represented as an ordered triple $(o_{i},\delta,o_{j})$, with $\delta\in\Delta$ denoting one of the following types:
  • Spatial co-occurrence: $o_{i}$ and $o_{j}$ must appear in the same frame, reflecting co-occurrence or proximity, e.g., for sentence “A person is standing beside a vase.” the relation will be (person, spatial, vase).
  • Attribute Dependency: $o_{i}$ and $o_{j}$ share a descriptive property such as color or size, capturing appearance-based attributes, e.g., for sentence “A person wears a black shirt.” the relation will be (person, attribute, black shirt).
  • Temporal Proximity: $o_{i}$ and $o_{j}$ occur in different frames within a short time interval, linking event progressions or changing states, e.g., for sentence “After a dog entered the room, a cat entered the room.” the relation will be (dog, temporal, cat).
  • Causal Order: $o_{i}$ and $o_{j}$ follow a clear temporal order indicating a cause–effect or prerequisite relation, e.g., for sentence “A little girl broke the vase.” the relation will be (little girl, causal, pieces).
These four types are chosen to comprehensively capture spatial co-occurrence, attribute dependency, temporal proximity, and cause-effect structures. As shown in Figure 1, we form the semantic-logical relation that effectively addresses a wide range of question-answering tasks. In particular, questions requiring temporal queries (“when does X happen?”), causal reasoning (“why did Y occur?”), attribute dependence (“What is the person wearing sunglasses doing?”) or detailed spatial constraints (“Who is standing next to the red car?”) can be more reliably answered by integrating these structured relations and contextual cues.
2.3 Iterative Semantic-Logical Temporal Search
Our algorithm iteratively searches for keyframes by semantic and logical reasoning, consisting of four main components: Frame Sampling, Object Detection and Scoring, Visual Semantic Logic Detection, and Distribution Update. The full pseudocode is provided in Algorithm 1.
1
2Function SemanticLogicalTemporalSearch($V,Q,K,\Delta_{t},\tau,\alpha,\gamma$)
3       $\mathcal{O},\mathcal{R}\leftarrow\text{ParseQuestion}(Q)$
4       // Extract key/cue objects and relations
5      $P\leftarrow\text{Uniform},B\leftarrow|V|,S\leftarrow\emptyset,N_{v}\leftarrow|V|$
6       // Initialize distribution and state
7      while $B>0$ and $|\mathcal{O}|>0$ do
8             $k\leftarrow\lfloor\sqrt{B}\rfloor$, $G\leftarrow\textnormal{{Grid}}(\textnormal{{Sample}}(P,k^{2}))$
9             // Adaptive grid sampling
10            $\Omega\leftarrow\textnormal{{DetectObjects}}(G)$
11             // Detect objects in sampled frames
12            foreach $t\in G$ do
13                   $C_{t}\leftarrow\textnormal{{CalculateBaseScore}}(\Omega[t])$
14                   // Base detection confidence
15                  foreach  $r_{type}\in\mathcal{R}$  do
16                         $\delta\leftarrow\textnormal{{Processrelation}}(r_{type},\Omega,\Delta_{t},\tau,\alpha,\gamma)$
17                         // relations require distinct processing
18                        $C_{t}\leftarrow C_{t}+\delta$
19                  $\textnormal{{UpdateScores}}(S,t,C_{t})$
20                   // Update global score registry
21
22            $\textnormal{{DiffuseScores}}(S,w)$
23             // Temporal context propagation
24            $P\leftarrow\textnormal{{NormalizeDistribution}}(S)$, $B\leftarrow B-k^{2}$
25             // Update sampling distribution
26            foreach $g\in\textnormal{{TopK}}(S,K)$ do
27                   if $\Omega[g]\cap\mathcal{O}\neq\emptyset$ then
28                         $\mathcal{O}\leftarrow\mathcal{O}\setminus\Omega[g]$
29                         // Remove identified key objects
30
31
32      return $\textnormal{{TopK}}(S,K)$
33       // Return top-K keyframes
Algorithm 1 Visual Semantic-Logical Search
2.3.1 Frame Sampling
To accelerate the search process, we do not examine each frame of the video exhaustively; instead, we adopt a distributed sampling strategy. Let $N_{v}$ denote the total number of frames in the video, and let $P$ be a sampling distribution initialized to uniform over all frames. We define the sampling process as:
$I_{s}=\mathrm{Sample}(P\odot N_{v},N_{s}),$ (1)
where $\mathrm{Sample}(\cdot,N_{s})$ selects a subset of $N_{s}$ frames according to the distribution $P\odot N_{v}$. To further accelerate subsequent steps, we stack the sampled frames into a $k\times k$ grid, which imposes a constraint on the sample size $N_{s}$. Specifically, we require:
$N_{s}\in\{k^{2}\mid k\in\mathbb{Z}\}\quad\text{and}\quad N_{s}<N_{v}.$ (2)
In practice, this ensures that the number of sampled frames can be reshaped into a compact 2D grid for efficient processing. Although $P$ is initially uniform, it can be adapted over multiple rounds of sampling to focus on frames of higher interest in the video.
2.3.2 Object Detection and Scoring
Method Training Required Searching Efficiency Overall Task Efficiency Matching Iteration TFLOPs $\downarrow$ Latency (sec) $\downarrow$ Latency (sec) $\downarrow$ Acc $\uparrow$ Static Frame Sampling Uniform-8 [38] Training-Based N/A N/A N/A 0.2 3.8 53.7 Dense Retrieval VideoAgent [10] Training-Based CLIP-1B [23] 840 536.5 30.2 34.9 49.2 T*-Retrieval [39] Training-Based YOLO-World-110M 840 216.1 28.6 32.2 57.3 Temporal Search T*-Attention [39] Training-Based N/A N/A 88.9 13.7 17.3 59.3 T*-Detector [39] Training-Free YOLO-World-110M 43 31.7 7.3 11.1 59.8 VSLS (ours)-Detector Training-Free YOLO-World-110M 49 33.3 7.8 11.6 61.5
Table 1: Evaluation of computational performance metrics across the LV-Haystack benchmark, presenting both search efficiency and end-to-end processing overhead (combining search and inference stages).
In this stage, we construct the detection search space by taking the union of both Key Objects and Cue Objects. For each iteration, we detect objects on the $N_{s}$ sampled frames using a lightweight model like YOLO-World [6] for high efficiency and score the frames based on detection confidence. Specifically, let $D_{t}$ be the set of detected objects in the frame at time $t$, $c_{o}$ the confidence of each detected object, and $w_{o}$ the corresponding weight. We define the frame score as:
$C_{t}=\max_{o\in D_{t}}\bigl{(}c_{o}\cdot w_{o}\bigr{)}.$ (3)
If the confidence score of any key object exceeds a predefined threshold, it is added to a list, thereby maintaining a record of frames where crucial targets have been identified for subsequent processing.
2.3.3 Visual Semantic Logic Detection
Beyond individual object detection and frame-level scoring, we further refine each frame’s confidence score by modeling higher-order logic relations among objects. Let $\mathcal{R}$ be a set of relations, where each relation $r\in\mathcal{R}$ is defined on a pair of objects $(o_{1},o_{2})$ and labeled with a relation type $r_{\text{type}}$. Denote by $C_{t}$ the confidence score for the frame at time $t$, and let $\alpha$ be a global scaling factor shared by all relations. We introduce an additional relation-specific weight $\gamma_{r_{\text{type}}}$ to control how strongly each type of logic affects the frame score. Formally, the updated frame confidence $C_{t}^{(r)}$ after satisfying relation $r$ is given by:
$C_{t}^{(r)}=C_{t}+\alpha\cdot\gamma_{r_{\text{type}}}.$ (4)
Spatial Relation. A spatial relation enforces that two objects $o_{1}$ and $o_{2}$ must co-occur in the same frame. Let $\Omega_{t}$ be the set of detected objects in frame $t$. If both $o_{1}\in\Omega_{t}$ and $o_{2}\in\Omega_{t}$, then the corresponding frame confidence is updated as:
$C_{t}\leftarrow C_{t}+\alpha\cdot\gamma_{\text{spatial}}.$ (5)
Attribute Relation. An attribute relation is satisfied when $o_{1}$ and $o_{2}$ share sufficient bounding-box overlap in the same frame. Let $\mathrm{overlap}$ be the ratio of their intersection area to the minimum of their individual bounding-box areas. If $\mathrm{overlap}>\tau$, we increment the frame confidence:
$C_{t}\leftarrow C_{t}+\alpha\cdot\gamma_{\text{attribute}}.$ (6)
Time Relation. A time relation checks whether two objects appear in temporally close frames. Suppose $t_{i}$ and $t_{j}$ ($t_{i}\leq t_{j}$) are sampled such that $|\,t_{j}-t_{i}|<\Delta_{t}$, where $\Delta_{t}$ is a threshold (e.g. five frames), if $o_{1}$ occurs in frame $t_{i}$ and $o_{2}$ in frame $t_{j}$, then both frames’ confidences are updated:
$C_{t_{i}}\leftarrow C_{t_{i}}+\alpha\cdot\gamma_{\text{time}},\quad C_{t_{j}}\leftarrow C_{t_{j}}+\alpha\cdot\gamma_{\text{time}}.$ (7)
Causal Relation. A causal relation models an ordering constraint, enforcing that $o_{1}$ must appear at an earlier time than $o_{2}$. Specifically, if $o_{1}\in\Omega_{t_{i}}$ and $o_{2}\in\Omega_{t_{j}}$ with $t_{i}<t_{j}$, we update the confidences of frames $t_{i}$ and $t_{j}$ by:
$C_{t_{i}}\leftarrow C_{t_{i}}+\alpha\cdot\gamma_{\text{causal}},\quad C_{t_{j}}\leftarrow C_{t_{j}}+\alpha\cdot\gamma_{\text{causal}}.$ (8)
Through these steps, frames that fulfill any relation criteria receive an incremental boost in confidence, thereby emphasizing contextually consistent detections. As we accumulate evidence from spatial, temporal, causal, and attribute relations, the system more accurately identifies the keyframes and interactions in the video.
2.4 Distribution Update
After each iteration of frame sampling, we merge newly obtained frame confidences into the global score distribution $\{S_{f}\}$ spanning all frames $f=1,2,\ldots,N_{v}$. When a frame $f$ is selected for detection, its score is assigned to the confidence value $C_{f}$, and the visitation counter $N_{v,f}$ is reset:
[…]
To incorporate temporal context, we diffuse this updated score to neighboring frames within a window of size $w$. Denoting each nearby index by $f\pm\delta$ (for $\delta\in[-w,w]$), we apply:
$S_{f\pm\delta}\leftarrow\max\left(S_{f\pm\delta},\,\frac{S_{f}}{1+|\delta|}\right).$ (10)
In this way, high-confidence frames raise the scores of close-by frames, reflecting temporal continuity. Following these local updates, the sampling distribution $P$ is refined using spline interpolation, and then normalized. This iteration proceeds until either the search budget $B$ is reached or all key objects have been successfully identified. The visualization of distribution in different iterations can be seen in Figure 3. Finally, the method outputs the top $K$ frames according to their terminal scores.
3 Experiment
3.1 Benchmark Datasets
The proposed VSLS is systematically evaluated across four benchmark datasets: a) LongVideoBench [38] for assessing long-context video-language comprehension capabilities; b) Video-MME [11] as the first comprehensive benchmark for multimodal video analytics; c) Haystack-LVBench, extended from LongVideoBench with human-annotated frame index answers; and d) Haystack-Ego4D, derived from Ego4D with similar annotations.
While LongVideoBench and Video-MME measure performance enhancement in QA accuracy, Haystack-Ego4D and Haystack-LVBench quantitatively evaluate keyframe selection accuracy through recall and precision metrics. Further details of these datasets, including descriptions, sizes, and usages are provided in Appendix C.
3.2 Evaluation Metrics
LongVideoBench Video-MME Model and Size Frame Video Length Model and Size Frame Video Length Long Medium Short Long Medium Short 900-3600s 180-600s 15-60s 30-60min 4-15min 0-2min GPT-4o [14] 8 47.1 49.4 67.3 GPT-4o 8 55.2 60.2 69.6 GPT-4o + T* 8 49.1 56.2 68.0 GPT-4o + T* 8 55.2 61.2 68.9 GPT-4o + VSLS (ours) 8 51.2 58.9 74 GPT-4o + VSLS (ours) 8 56.9 60.7 68.2 InternVL 2.5-78B [5] 8 55.7 57.3 74.0 InternVL 2.5-78B 8 52.6 55.5 55.9 InternVL 2.5-78B + VSLS (ours) 8 58.0 61.5 74.0 InternVL 2.5-78B + VSLS (ours) 8 57.7 57.5 59.0 GPT-4o 32 53.8 56.5 74.0 GPT-4o 32 55.2 61.0 71.4 GPT-4o + T* 32 55.3 58.8 72.0 GPT-4o + T* 32 55.2 61.6 72.6 GPT-4o + VSLS (ours) 32 54.2 60.0 76.0 GPT-4o + VSLS (ours) 32 55.2 61.9 71.9 LLaVA-OneVision-QWen2-78B-ov 32 59.3 63.9 77.4 LLaVA-OneVision-78B 32 60.0 62.2 66.3 PLLaVA-34B 32 49.1 50.8 66.8 VideoLLaMA 2 32 57.6 59.9 62.4 LLaVA-Video-78B-Qwen2 128 59.3 63.9 77.4 Oryx-1.5 128 59.3 65.3 67.3 mPLUG-Owl3-7B 128 53.9 58.8 73.7 Aria-8x3.5B 256 58.8 67.0 67.6 GPT-4o (0513) 256 61.6 66.7 76.8 Gemini-1.5-Pro (0615) 1/0.5 fps 67.4 74.3 75.0
Table 2: Downstream task evaluation results on two benchmarks. All the accuracy(%) in black is based on our replication results. We also cite the reported accuracy of the SOTA models in gray (note that their settings may differ from ours, and these results are not guaranteed to be reproducible) and list their number of frames used for QA inference, which we include for full transparency and openness.
3.2.1 Evaluation Metrics for Search Utility
Our assessment framework emphasizes both the effectiveness and efficiency of search functions. In terms of search effectiveness, we use three metrics to compare model-predicted keyframes with human annotations, considering both individual frames and entire sets. This approach tackles the issue that multiple valid keyframe sets might exist for a single query. For frame-to-frame comparisons, we evaluate the alignment between a model-predicted frame $f_{\text{pt}}$ and a human-annotated frame $f_{\text{gt}}$ from two perspectives:
Temporal coverage evaluates the coverage of ground truth frames by predicted frames in the temporal perspective, which can be described as:
$T_{\text{cover}}(T_{\text{pt}},T_{\text{gt}})=\frac{\sum\limits_{i=1}^{|N_{\text{gt}}|}\mathbb{I}\left[\min\limits_{j}\left|t_{\text{gt}}^{i}-t_{\text{pt}}^{j}\right|\leq\delta\right]}{|N_{\text{gt}}|},$ (11)
where $T_{\text{pt}}$ represents the set of predicted timestamps and $T_{\text{gt}}$ represents the set of ground truth timestamps. In this formula, $|N_{\text{gt}}|$ denotes the total number of ground truth frames, $t_{\text{gt}}^{i}$ is the timestamp of the $i$-th ground truth frame, $t_{\text{pt}}^{j}$ is the timestamp of the $j$-th predicted frame, $\delta$ is the temporal similarity threshold defining the maximum acceptable time difference, and $\mathbb{I}[\cdot]$ is the indicator function that returns 1 when the condition is satisfied and 0 otherwise.
Visual Similarity uses the Structural Similarity Index Measure (SSIM) [1] to evaluate the visual resemblance between $f_{\text{pt}}$ and $f_{\text{gt}}$ based on structural details, luminance, and contrast. For set-to-set comparisons, the main challenge is defining the similarity between two sets. We use Precision $P$ and Recall $R$ as complementary metrics. Precision checks if each model-selected frame matches at least one reference frame, while Recall ensures that reference frames are adequately represented in the model’s output. Given the ground truth keyframe set $F_{\text{gt}}=\{f^{j}_{\text{gt}}\}^{n}_{j=1}$ and the predicted keyframe set $F_{\text{pt}}=\{f^{j}_{\text{pt}}\}^{m}_{i=1}$, we formally define the multimodal retrieval quality metrics as follows:
$P(F_{\text{pt}},F_{\text{gt}})$ $=\frac{1}{|F_{\text{pt}}|}\sum_{f^{i}_{\text{pt}}\in F_{\text{pt}}}\max_{f^{j}_{\text{gt}}\in F_{\text{gt}}}\phi(f^{i}_{\text{pt}},f^{j}_{\text{gt}});$ (12a)
$R(F_{\text{pt}},F_{\text{gt}})$ $=\frac{1}{|F_{\text{gt}}|}\sum_{f^{j}_{\text{gt}}\in F_{\text{gt}}}\max_{f^{i}_{\text{pt}}\in F_{\text{pt}}}\phi(f^{j}_{\text{gt}},f^{i}_{\text{pt}}),$ (12b)
where $\phi(\cdot,\cdot)$ represents an extensible multimodal similarity metric function.
Figure 3: Sample weight evolution under VSLS optimization for keyframe selection. Top: 16 iterations show progressive convergence toward Ground Truth (red). Bottom: 15 iterations demonstrate similar alignment. Yellow highlights indicate precise matches between algorithm outputs (green) and manual annotations.
3.2.2 Evaluation Metrics for Search efficiency
Existing studies [10, 21, 31, 33, 36] have mainly concentrated on optimizing task-specific performance metrics while neglecting computational efficiency in temporal search operations. To systematically analyze this dimension, our evaluation framework incorporates two criteria: 1) FLOPs representing arithmetic operation complexity, and 2) Latency recording real-world execution duration.
3.3 Evaluation of Search Framework efficiency
Current approaches for keyframe selection can be broadly categorized into three paradigms: statistic-based frame sampling, dense feature retrieval-based selection, and temporal search-based methods. As shown in Table 1, while uniform sampling achieves the fastest processing speed, its ignorance of frame semantics severely limits downstream task effectiveness. Although dense feature retrieval methods attain moderate accuracy improvements (57.3%), their exhaustive frame processing demands 4.2$\times$ more TFLOPs and introduces 4.5$\times$ higher latency than our temporal search approach. Crucially, our method introduces four visual semantic logic detectors during temporal search while maintaining comparable execution time to T* methods. This strategic design elevates downstream task accuracy to 61.5%, achieving the best performance-efficiency trade-off.
3.4 Visual Semantic Logic Search Performance
As demonstrated in Table 3, we evaluate VSLS on LongVideoBench from two critical perspectives: visual similarity (measured by precision and recall) and temporal coverage. Our method achieves state-of-the-art performance across all metrics. Specifically, under the 32-frame setting, VSLS attains a precision of 74.5% and recall of 92.5%, outperforming all baselines in visual similarity. More notably, the temporal coverage of VSLS reaches 41.4%, surpassing the second-best method (T* at 36.5%) by 13.4%—the largest margin among all comparisons. This significant improvement highlights the effectiveness of our visual semantic logic detection modules in identifying query-relevant keyframes with both semantic alignment and temporal completeness.
These results empirically prove the foundational hypothesis that extracting semantic and logical cues from textual queries facilitates precise detection of query-relevant video frames. The marked improvements in visual similarity and temporal coverage demonstrate that VSLS successfully captures corresponding keyframes while maintaining temporal coherence through the visual-logical alignment.
Method Frame LongVideoBench Precision $\uparrow$ Recall $\uparrow$ Time $\uparrow$ Static Frame Sampling Method Uniform [38] 8 56.0 72.0 6.3 Uniform 8 60.7 80.4 4.7 Uniform 32 58.7 81.6 24.9 Uniform 32 60.2 85.0 8.1 Dense Retrieval Method VideoAgent [10] 10.1 58.8 73.2 8.5 Retrieval-based [39] 8 63.1 65.5 6.3 Retrieval-based 32 59.9 80.8 21.8 Temporal Searching Method T* [39] 8 58.4 72.7 7.1 T* 8 75.3 88.2 26.2 VSLS (ours) 8 75.6 88.6 26.3 T* 32 58.3 83.2 28.2 T* 32 74.0 90.3 36.5 VSLS (ours) 32 74.5 92.5 41.4
Table 3: Results of searching utility on LongVideoBench. 8-frame setting bests are underlined, and 32-frame setting bests are in bold. Gray font indicates results from the original paper.
3.5 Downstream Video QA Performance
To further demonstrate the advantages of our VSLS, we conduct downstream video QA performance experiments on LongVideoBench and Video-MME datasets. As shown in Table 2, we categorize the video length into three groups, namely Long, Medium, and Short, ranging from 15 to 3600s, and 0 to 60 mins respectively. Our method consistently achieves the highest accuracy in the long video category across different numbers of frames and downstream QA models. Compared to our baseline method T*, we observe significant performance improvements by incorporating our selected visual semantic logic relation, which is shown in Figure 1. These results validate that our approach can capture and utilize visual logical relations in long video content, leading to superior video QA performance.
4 Analysis
4.1 Time Complexity
The proposed framework can be divided into two stages. In the initial phase, VLMs such as LlaVA-7B and GPT-4o, are employed to extract a semantic set $\mathcal{S}$ from a video $V$ containing $n$ frames. This semantic set $\mathcal{S}$ comprises three semantic components: target objects, cue objects, and their relational information. Notably, the size of $\mathcal{S}$ is constrained by our prompt design. Following this, the subsequent phase focuses on key-frame identification and operates as a heuristic search algorithm. Specifically, this process iteratively selects $k$ candidates from the $n$ frames through a scoring mechanism guided by an evaluation function $h(\cdot,\mathcal{S})$. Furthermore, the score distribution $\text{scores}[n]$ is dynamically updated based on the YOLO-World detector’s output, ensuring continuous refinement of the selection process.
Our analysis will focus on the number of YOLO-World detections required during the search process, which is the primary computational bottleneck of our framework. This is because each detection involves running a deep neural network, making it the most resource-intensive step. Reducing the number of detections directly improves efficiency without compromising accuracy. In each iteration, the YOLO-World detector is leveraged to match the selected $k$ frames with the objects and relations in $\mathcal{S}$, resulting in $k$ YOLO-World detections. The search terminates when all objects are detected and their relations are satisfied or when the number of iterations reaches the predefined search budget, defined as $\mathrm{min}(1000,0.1\times V_{t})$, where $V_{t}$ denotes the duration of the video $V$ in seconds. In the worst cases, for a video longer than 10,000 frames where each iteration fails to identify any object and relation, the number of iterations will be clamped at 1,000. Under ideal conditions, the evaluation function $h(\cdot,\mathcal{S})$ consistently identifies all target frames and leads to the optimal score distribution; the algorithm is similar to selecting top-$k$ candidates from n frames within $\mathcal{O}(|\mathcal{S}|\log n)$ iterations [39]. Therefore, the average time complexity of our algorithm is $\mathcal{O}(|\mathcal{S}|k\log n)$ YOLO-World inferences.
Experimental results also demonstrate that integrating relational information into the search algorithm incurs negligible computational overhead compared to the baseline T* approach. On the LV-Haystack benchmark, the average iteration count increases from 42.94 (T*) to 48.82 iterations, representing a modest 13.69% rise in the time cost.
Figure 4: Distribution of semantic-logical relations identified by VSLS in LongVideoBench and Video-MME datasets. Spatial relations dominate both (57.7% vs. 72.9%), with notable differences in causal relations (33.7% vs. 19.2%) and attribute relations (4.7% vs. 3.1%). Time relations remain least frequent (3.8% vs. 4.9%).
4.2 Ablation Study of Four Relations
Figure 4 illustrates the distribution of four logic relation types across LongVideoBench and Video-MME datasets, where spatial relations predominate, followed by attribute relations. In Table 4, we extract samples containing different relation types from LongVideoBench to compare the object detection-based T* method with our VSLS approach. Experimental results demonstrate that VSLS achieves significant improvements across both image similarity metrics (SSIM Precision and SSIM Recall). Additionally, temporal coverage shows marked enhancement for attribute, spatial, and causal relations, with spatial relations exhibiting the most substantial improvement (21.3% increase over T*). For the time relation category, we observe a slight decrease in temporal coverage, which may be attributed to the relative scarcity of time relation samples in the dataset, limiting the opportunity to demonstrate the advantages of VSLS. Nevertheless, Figure 1 provides visual evidence of how effectively leveraging time relations can facilitate downstream question-answering tasks.
LongVideoBench
Logic Type Method Precision $\uparrow$ Recall $\uparrow$ TC $\uparrow$
T* 72.9 88.7 37.5
Spatial VSLS (ours) 73.6 91.4 45.5
T* 71.8 87.6 38.5
Attribute VSLS (ours) 72.7 90.9 42.1
T* 76.7 89.2 37.3
Time VSLS (ours) 77.5 92.5 36.1
T* 74.7 92.4 38.6
Casual VSLS (ours) 74.7 93.8 39.6
Table 4: Results of searching utility on LongVideoBench. Precision: SSIM Precision; Recall: SSIM Recall; TC: Temporal Coverage. Results demonstrate significant improvements in two image-similarity metrics across all four categories of samples.
5 Related Work
Challenges in Long Video Understanding: Long video understanding is inherently more challenging than short-video or image-based tasks due to its rich temporal dynamics and massive redundancy [22, 43, 42]. The large number of frames increases both memory and computational requirements, making straightforward dense sampling infeasible. Moreover, crucial events may span distant timestamps, demanding high-capacity models to capture long-range dependencies [25, 26, 3, 34]. Meanwhile, the diverse and continuous visual content raises noise and distractors; thus, strategies to effectively locate or distill essential parts of the video are of primary importance [44, 7, 37, 39].
Existing Solutions based on VLMs typically share three core ideas: 1) video sampling or retrieval for efficiency, 2) multi-stage or interactive reasoning to handle complex questions, and 3) compact representation to accommodate the VLM’s limited context window. For instance, retrieval-based pipelines partition a video into segments and employ a learned or rule-based retriever to identify the relevant chunks before passing them to a VLM [20, 8, 9, 13]. Other lines of research compress each frame into minimal tokens to reduce computational overhead [16, 2, 27], or adopt a streaming mechanism to propagate memory representations along the temporal axis [22, 35, 19]. Beyond these efficiency-oriented approaches, LLM/VLM-as-planner frameworks factorize the process into a series of perception queries, enabling an agent to fetch additional frame-level details if needed [31, 45, 18].
6 Conclusion
In this paper, we present Visual Semantic-Logical Search (VSLS), a novel framework that efficiently selects semantically keyframes for long video understanding by decomposing logical relationships between textual queries and visual elements. VSLS based on four defined logical dependencies (spatial co-occurrence, temporal proximity, attribute dependency, and causal order), significantly outperforms existing methods while sampling only 1.4% of video frames. The 8.7% improvement in GPT-4o’s long video QA accuracy demonstrates that query-guided visual semantic logic search effectively bridges the gap between textual queries and visual content. VSLS’s plug-and-play nature enables seamless integration with existing pipelines, making it practical for real-world applications. Future work could consider more logical relations, learnable search methods, enhancing interpretability, and exploring more downstream tasks.
Supplementary Material
Part I Appendix
[…]
Appendix A Performance
Long-form video understanding presents unique challenges due to the complexity of temporal dynamics and cross-modal interactions in extended durations (900-3,600 seconds). Our comprehensive evaluation of the LVB-XL benchmark reveals significant performance gaps between existing approaches. While large-scale models like GPT-4o (32 frames) and InternVL 2.5-78B (16 frames) have demonstrated competence in short-video tasks, their direct application to long-form content (marked by circle sizes proportional to model parameters) yields suboptimal results (53.8% and 56.5% accuracy respectively).
Our Visual Semantic-Logical Search (VSLS) framework addresses these limitations through systematically defined four fundamental logical dependencies. This advancement enables consistent performance improvements across different architecture scales, elevating GPT-4o to 54.2% (+0.4pp) and achieving a remarkable 62.4% (+5.9pp) for InternVL 2.5-78B on this benchmark. The comparative analysis further suggests that VSLS’s gains become particularly pronounced when processing longer visual sequences, highlighting its effectiveness in modeling extended temporal contexts.
Appendix B Analysis of the Impact of Search Frame Count
Figure 5: Performance improvement with increasing search frames. VSLS consistently enhances accuracy and reaches near-human oracle performance at 64 frames.
This section investigates the impact of the number of search frames on the performance of our Visual Language Models (VLMs) in the context of LongVideoBench.
Figure 5 in the T* framework study empirically demonstrates the non-monotonic relationship between input frame quantity and model accuracy on the LongVideoBench XL benchmark. Through systematic experimentation across 18 state-of-the-art VLMs, this visualization reveals a critical phenomenon: excessive frame inputs degrade performance for models lacking temporal redundancy mitigation mechanisms.
Appendix C Details of Datasets
C.1 Details of Video-MME
The Video-MME (Video Multi-Modal Evaluation) dataset represents the first comprehensive benchmark tailored to assess the capabilities of Vision-Language Models (VLMs) in video understanding. Aiming to address limitations in existing benchmarks, it emphasizes diversity, temporal complexity, and multi-modal integration while ensuring high-quality human annotations. The dataset contains 900 carefully curated videos across six primary domains—Knowledge, Film and Television, Sports Competition, Artistic Performance, Life Record, and Multilingual—with 30 fine-grained subcategories such as astronomy, esports, and documentaries. These videos vary significantly in duration, ranging from short clips (11 seconds) to long-form content (up to 1 hour), enabling robust evaluation across temporal scales.
Each video is paired with expert-annotated multiple-choice questions (2,700 QA pairs in total), rigorously validated to ensure clarity and reliance on visual or multi-modal context. Questions span 12 task types, including action recognition, temporal reasoning, and domain-specific knowledge, with a focus on scenarios where answers cannot be inferred from text alone. To quantify temporal complexity, the dataset introduces certificate length analysis, revealing that answering questions often requires understanding extended video segments (e.g., median lengths of 26 seconds for short videos and 890.7 seconds for long videos), surpassing the demands of prior benchmarks like EgoSchema.
Video-MME serves as a universal benchmark, applicable to both image- and video-focused MLLMs, and exposes key challenges for future research. These include improving architectures for long-sequence processing, developing datasets for complex temporal reasoning, and enhancing cross-modal alignment. By providing a rigorous evaluation framework, Video-MME aims to drive progress toward MLLMs capable of understanding dynamic, real-world scenarios.
C.2 Details of LongVideoBench
The LongVideoBench benchmark pioneers the evaluation of long-context interleaved video-language understanding in VLMs, addressing critical gaps in existing benchmarks through its focus on detailed retrieval and temporal reasoning over hour-long multimodal inputs. Designed to overcome the “single-frame bias” prevalent in prior video benchmarks, the novel referring reasoning paradigm enables models to locate and analyze specific contexts within extended sequences. The data set comprises 3,763 web-sourced videos that span various themes - movies, news, life vlogs, and knowledge domains (including art, history, and STEM) - with durations progressively grouped into four levels: 8-15 seconds, 15-60 seconds, 3-10 minutes, and 15-60 minutes. Each video is paired with aligned subtitles, forming interleaved multimodal inputs that mimic real-world viewing scenarios.
The benchmark features 6,678 human-annotated multiple-choice questions categorized into 17 fine-grained task types across two levels: Perception (requiring object/attribute recognition in single scenes) and Relation (demanding temporal/causal reasoning across multiple scenes). Questions incorporate explicit referring queries (e.g., “When the woman descends the rocky hill…”) that anchor reasoning to specific video moments, with an average question length of 43.5 words to ensure precision. Temporal complexity is quantified through duration-grouped analysis, where models must process up to 256 frames (at 1 fps) for hour-long videos, significantly exceeding the demands of predecessors like EgoSchema (180s videos).
C.3 Details of LV-Haystack
The LV-Haystack benchmark establishes the first comprehensive evaluation framework for temporal search in long-form video understanding, addressing critical limitations in existing synthetic needle-in-haystack benchmarks through real-world video annotations and multi-dimensional evaluation metrics. Designed to assess models’ ability to locate minimal keyframe sets (typically 1-5 frames) from hour-long videos containing tens of thousands of frames, the dataset comprises 3,874 human-annotated instances spanning 150 hours of video content across two distinct categories: egocentric videos from Ego4D (101 hours) and allocentric videos from LongVideoBench (57.7 hours).
Organized into Haystack-Ego4D and Haystack-LVBench subsets, the benchmark features videos averaging 24.8 minutes in length (max 60 minutes) with 44,717 frames per video. Each instance contains:
  • Expert-curated multi-choice questions requiring temporal reasoning (15.9 questions/video);
  • Human-annotated keyframe sets (4.7 frames/question for egocentric, 1.8 frames/question for allocentric);
  • Temporal and visual similarity metrics for precise search evaluation.
C.4 Details of Ego-4D
1
2Function SemanticLogicalTemporalSearch($V,Q,K,\Delta_{t},\tau,\alpha,\gamma$):
       $\mathcal{O},\mathcal{R}\leftarrow\text{ParseQuestion}(Q)$ ;
        // Extract key/cue objects and relationships
       $P\leftarrow\text{Uniform},B\leftarrow|V|,S\leftarrow\emptyset,N_{v}\leftarrow|V|$ ;
        // Initialize distribution and state
3
4      while $B>0$ and $|\mathcal{O}|>0$ do
             $k\leftarrow\lfloor\sqrt{B}\rfloor$, $G\leftarrow\textnormal{{Grid}}(\textnormal{{Sample}}(P,k^{2}))$ ;
              // Adaptive grid sampling
             $\Omega\leftarrow\textnormal{{DetectObjects}}(G)$ ;
              // Detect objects in sampled frames
5
6            foreach $g\in G$ do
                   $C_{g}\leftarrow\textnormal{{CalculateBaseScore}}(\Omega[g])$ ;
                    // Base detection confidence
7                   foreach $r\in\mathcal{R}$ do
8                         if $r.\text{type}=\text{Spatial}$ then
9                               $C_{g}\leftarrow C_{g}+\alpha\gamma_{\text{spatial}}\cdot\textnormal{{CheckSpatialRelationship}}(r,\Omega[g])$
10                         else if $r.\text{type}=\text{Temporal}$ then
11                               $C_{g}\leftarrow C_{g}+\alpha\gamma_{\text{time}}\cdot\textnormal{{CheckTemporalRelationship}}(r,\Omega,\Delta_{t})$
12                         else if $r.\text{type}=\text{Causal}$ then
13                               $C_{g}\leftarrow C_{g}+\alpha\gamma_{\text{causal}}\cdot\textnormal{{CheckCausalRelationship}}(r,\Omega)$
14                         else if $r.\text{type}=\text{Attribute}$ then
15                               $C_{g}\leftarrow C_{g}+\alpha\gamma_{\text{attr}}\cdot\textnormal{{CheckAttributeRelationship}}(r,\Omega[g],\tau)$
16
                  $\textnormal{{UpdateScores}}(S,g,C_{g})$ ;
                    // Update global score registry
17
18
            $\textnormal{{DiffuseScores}}(S,w)$ ;
              // Temporal context propagation
             $P\leftarrow\textnormal{{NormalizeDistribution}}(S)$, $B\leftarrow B-k^{2}$ ;
              // Update sampling distribution
19
20            foreach $g\in\textnormal{{TopK}}(S,K)$ do
21                   if $\Omega[g]\cap\mathcal{O}\neq\emptyset$ then
                         $\mathcal{O}\leftarrow\mathcal{O}\setminus\Omega[g]$;
                          // Remove identified key objects
22
23
24
      return $\textnormal{{TopK}}(S,K)$ ;
        // Return top-$k$ keyframes
25
Algorithm 2 The completed Visual Semantic-Logical Search
The Ego4D (Egocentric Computer Vision Benchmark) dataset establishes a transformative foundation for advancing research in first-person visual perception through unprecedented scale, diversity, and multi-modal integration. Designed to overcome limitations in existing egocentric datasets, it captures 3,670 hours of unscripted daily activities from 931 participants across 74 global locations and 9 countries, spanning household, workplace, leisure, and outdoor scenarios. The dataset features 30+ fine-grained activity categories including carpentry, social gaming, and meal preparation, with videos ranging from brief interactions (8-minute clips) to extended continuous recordings (up to 10 hours), enabling comprehensive analysis of long-term behavioral patterns.
Each video is enriched with multi-modal annotations totaling 3.85 million dense textual narrations (13.2 sentences/minute), coupled with 3D environment meshes, eye gaze tracking, stereo vision, and synchronized multi-camera views. Rigorous privacy protocols ensure ethical data collection, with 612 hours containing unblurred faces/audio for social interaction studies. The benchmark suite introduces five core tasks organized across temporal dimensions:
  • Episodic Memory: Temporal localization of natural language queries (74K instances) and 3D object tracking using Matterport scans;
  • Hand-Object Interaction: State change detection (1.3M annotations) with PNR (point-of-no-return) temporal localization;
  • Social Understanding: Audio-visual diarisation (2,535h audio) and gaze-directed communication analysis;
  • Action Forecasting: Anticipation of locomotion trajectories and object interactions.
Quantitative analysis reveals the dataset’s complexity: hand-object interactions involve 1,772 unique verbs and 4,336 nouns, while social scenarios contain 6.8 participant interactions per minute on average. Multi-modal fusion experiments demonstrate performance gains, with 3D environment context improving object localization accuracy by 18.7% compared to RGB-only baselines. State-of-the-art models achieve 68.9% accuracy in action anticipation tasks, yet struggle with long-term forecasting (41.2% accuracy for 5s predictions), highlighting critical challenges in temporal reasoning.
Ego4D’s unique integration of egocentric video with complementary modalities (IMU data in 836h, gaze tracking in 45h) enables novel research directions in embodied AI and augmented reality. The dataset exposes fundamental limitations in current architectures, particularly in processing hour-long video contexts and synthesizing cross-modal signals—only 23% of tested models effectively utilized audio-visual synchronization cues. By providing standardized evaluation protocols and curated challenge subsets, Ego4D serves as a universal testbed for developing perceptive systems capable of understanding persistent 3D environments and complex human behaviors.
Appendix D Visual Semantic-Logical Search Algorithm
The Visual Semantic-Logical Search algorithm (Algorithm 2) represents a sophisticated approach to searching visual content based on semantic understanding, logical relationships, and temporal context.
D.1 Algorithm Overview and Core Components
The algorithm operates as an adaptive search framework that intelligently explores video content (represented as set $V$) to locate frames matching semantic-logical query requirements ($Q$). Unlike traditional linear search methods, it employs a probabilistic sampling strategy that dynamically adjusts based on confidence scores from multiple relationship types.
D.1.1 Initialization Phase
The process begins by parsing the input query $Q$ into two fundamental components:
  • $\mathcal{O}$: A set of key objects or entities to identify;
  • $\mathcal{R}$: A collection of relationships (spatial, temporal, causal, and attribute) that must be satisfied.
The algorithm initializes with a uniform probability distribution ($P$) across all video frames, establishing a budget ($B$) equivalent to the total number of frames ($|V|$), and creating an empty score registry ($S$) to track confidence values. This approach ensures unbiased initial exploration before evidence-guided refinement.
D.1.2 Adaptive Sampling Strategy
Rather than exhaustively processing every frame, the algorithm employs a square-root scaling sampling strategy where $k=\lfloor\sqrt{B}\rfloor$ determines the sampling density. This provides a mathematical balance between exploration breadth and computational efficiency. The Grid function organizes sampled frames into a structured representation that preserves spatial-temporal relationships, facilitating subsequent relationship analysis.
D.1.3 Multi-modal Object Detection
The DetectObjects function applies state-of-the-art computer vision techniques to identify objects within each sampled frame. This step leverages deep neural networks pre-trained on diverse visual datasets, enabling recognition of a wide range of entities with their corresponding confidence scores and spatial locations within frames.
D.1.4 Score Propagation and Distribution Update
The DiffuseScores function implements a temporal context propagation mechanism that spreads confidence values to neighboring frames, acknowledging that relevant content likely extends beyond individual frames. This diffusion creates a smoothed confidence landscape that guides subsequent sampling.
After each iteration, the algorithm normalizes the accumulated scores to form an updated probability distribution, focusing future sampling on promising regions while maintaining exploration potential in unexamined areas.
D.1.5 Convergence Criteria and Termination
The search continues until either:
  1. 1.
    The sampling budget ($B$) is exhausted, indicating comprehensive coverage of the video content;
  2. 2.
    All target objects ($\mathcal{O}$) have been successfully identified at satisfactory confidence levels.
This dual-termination approach balances thoroughness with efficiency, preventing unnecessary computation once objectives are met.
D.1.6 Result Generation
The algorithm concludes by returning the top-$k$ frames with highest confidence scores, representing the most relevant video segments that satisfy the semantic-logical query requirements. These keyframes provide a concise summary of the content matching the user’s information needs.
D.2 Implementation Considerations
The algorithm’s performance depends on several configurable parameters:
  • $\Delta_{t}$: Temporal window size for relationship analysis;
  • $\tau$: Confidence threshold for attribute matching;
  • $\alpha$: Global relationship influence factor;
  • $\gamma$: Type-specific relationship weights.
These parameters can be tuned based on application requirements, video characteristics, and computational constraints. The algorithm’s modular design allows for straightforward substitution of specific component implementations (e.g., different object detectors or relationship checkers) without altering the overall framework.
D.3 Computational Complexity Analysis
The time complexity scales with $O(\sqrt{N})$ where $N$ is the total number of frames, significantly improving upon linear approaches. Space complexity remains $O(N)$ to maintain the probability distribution and score registry. The algorithm intelligently balances exploration and exploitation through its adaptive sampling approach, making it particularly suitable for large-scale video analysis tasks where exhaustive processing would be prohibitive.
D.4 Technical Implementation Details
D.4.1 Object Detection and Feature Extraction
To achieve real-time performance, the object detection module utilizes pre-trained deep convolutional neural network architectures, particularly variants based on Fast R-CNN and YOLO series. The system employs a two-stage detection strategy:
  • Preliminary Detection: Using lightweight models to rapidly identify potential regions;
  • Fine-grained Classification: Applying more sophisticated models for detailed classification on high-confidence regions.
The feature extraction process leverages self-attention mechanisms from Visual Transformers (ViT), generating rich semantic embeddings robust to various visual variations such as scale, rotation, and illumination. Each identified object is associated with a feature vector $f_{i}\in\mathbb{R}^{d}$, where $d=512$ represents the dimensionality of the embedding space.
D.4.2 Mathematical Formulations for Relationship Assessment
The evaluation of various relationship types is based on precise mathematical definitions:
Spatial Relationships
Given bounding boxes $B_{i}=(x_{i},y_{i},w_{i},h_{i})$ and $B_{j}=(x_{j},y_{j},w_{j},h_{j})$ for two objects, the confidence for a spatial relationship $r_{spatial}$ is calculated as:
$C_{\text{spatial}}(B_{i},B_{j},r)=\phi_{r}(B_{i},B_{j})\cdot\psi(B_{i})\cdot\psi(B_{j}),$ (13)
where $\phi_{r}$ is a relationship-specific compatibility function and $\psi$ is the object detection confidence. For example, the compatibility for a “contains” relationship is defined as:
$\phi_{\text{contains}}(B_{i},B_{j})=\frac{\text{IoU}(B_{i},B_{j})}{\text{Area}(B_{j})}.$ (14)
Temporal Relationships
Temporal relationships are calculated by evaluating object behavior patterns across a sequence of frames $\{F_{t},F_{t+1},...,F_{t+\Delta_{t}}\}$:
$C_{\text{temporal}}(O_{i},O_{j},r,\Delta_{t})=\prod_{k=0}^{\Delta_{t}-1}T_{r}(O_{i}^{t+k},O_{j}^{t+k+1}),$ (15)
where $T_{r}$ is a relationship-specific temporal transition matrix and $O_{i}^{t}$ represents the state of object $i$ at time $t$.
Causal Relationships
Causal relationships utilize a Bayesian network framework to compute conditional probabilities:
$C_{\text{causal}}(E_{i},E_{j})=P(E_{j}|E_{i})\cdot\log\frac{P(E_{j}|E_{i})}{P(E_{j})},$ (16)
where $E_{i}$ and $E_{j}$ represent the presumed cause event and effect event, respectively.
Attribute Relationships
Attribute evaluation employs cosine similarity metrics between feature vectors and attribute prototypes:
$C_{\text{attr}}(O_{i},a)=\max(0,\cos(f_{i},p_{a})-\tau),$ (17)
where $p_{a}$ is the prototype vector for attribute $a$ and $\tau$ is the minimum similarity threshold.
D.4.3 Score Propagation Algorithm
Temporal score propagation is implemented through a weighted diffusion process, analogous to heat diffusion on a graph structure:
$S^{\prime}(t)=S(t)+\sum_{k\in\mathcal{N}(t)}w_{k,t}\cdot S(k),$ (18)
where $\mathcal{N}(t)$ represents the temporal neighborhood of frame $t$, and $w_{k,t}$ is a weight based on temporal distance, defined as:
$w_{k,t}=\exp\left(-\frac{|k-t|^{2}}{2\sigma^{2}}\right),$ (19)
where $\sigma$ controls the diffusion range.
D.4.4 Adaptive Sampling Optimization
The sampling strategy is further improved through a dynamically adjusted Thompson sampling method, modeling the probability distribution $P$ as a Beta distribution with shape parameters updated through previous observations:
$P(t)\sim\text{Beta}(\alpha_{t}+\sum_{i}S_{i}(t),\beta_{t}+n-\sum_{i}S_{i}(t)),$ (20)
where $\alpha_{t}$ and $\beta_{t}$ are prior hyperparameters and $n$ is the total number of observations.
D.5 Practical Application Examples
In practical visual search scenarios, the algorithm processes complex queries such as “a person wearing a blue shirt sits down at a table and then picks up a coffee cup”:
  • Query parsing identifies key objects (person, shirt, table, coffee cup) and relationships (blue attribute, sitting action, temporal before-after relation, spatial proximity);
  • Adaptive sampling selects representative frames from the video;
  • Multi-relationship evaluation integrates various sources of evidence;
  • Score propagation establishes a unified confidence landscape across related frame sets;
  • Result generation provides a concise summary of the most relevant segments in the video.
This semantic-logical-temporal search framework represents a significant advancement in multimodal content retrieval, enabling natural language queries that incorporate complex relationships across objects, time, and causal chains.
Appendix E Case Study of VSLS Keyframe Selection
Figure 6: Qualitative comparison of frame selection strategies demonstrates VSLS’s ability to pinpoint query-critical moments (e.g., the subject presenting pink objects) with temporal precision, while baseline approaches exhibit color misinterpretation (brown) due to suboptimal frame choices. VSLS maintains superior temporal diversity and content relevance, effectively avoiding the redundant selections observed in comparative methods.
As shown in Figure 6, the VSLS framework demonstrates its effectiveness through a video question-answering case study involving temporal handwriting analysis. The experiment focuses on distinguishing between two sequential events: a brown pen writing “guitar” at 2 seconds and a pink pen rewriting the same word at 3 seconds, with the query requiring identification of the second occurrence’s pen color.
VSLS’s analytical process unfolds through three interpretable phases:
  • Semantic Logic Extraction: Identifies core visual entities (handwritten text, pen, paper) and constructs temporal relationships through triplet formulation: (text, time, pen), establishing the framework for tracking writing instrument changes;
  • Temporal Relevance Scoring: The gray relevance curve reveals precise temporal localization, with peak scores aligning perfectly with ground truth positions at 2s and 3s, contrasting sharply with baseline methods’ random fluctuations;
  • Search Pattern Visualization: Demonstrates VSLS’s focused inspection near critical moments versus uniform sampling’s scattered temporal coverage, explaining the baseline’s failure to detect the pink pen.
This case study yields two critical insights about VSLS’s temporal reasoning:
  1. 1.
    Sequential Event Disambiguation: The system successfully differentiates between near-identical visual events through:
    • First writing instance: Brown pen detection (false positive);
    • Second writing instance: Pink pen detection (true positive).
  2. 2.
    Explanation of answer generation disparity: VSLS produces the correct answer (“Pink”) versus uniform sampling’s erroneous baseline (“Brown”) due to temporal reasoning failures.
The spatial-temporal alignment between relevance peaks and ground truth positions confirms VSLS’s unique capacity to synchronize semantic logic with visual evidence flow. This case particularly highlights the method’s superiority in scenarios requiring precise discrimination of recurrent events with subtle visual variations.
Appendix F Iteration Analysis
Figure 7: The comparative visualization of iteration counts on the medium-length video subset of the Video-MME dataset demonstrates that our method consistently requires a higher number of iterations compared to the T* approach.
As shown in Fig 7, incorporating relations into the search algorithm will increase the average number of iterations for the video of medium length in the Video-MME dataset from 15.9 to 23.8. The overall distribution of video iteration will not be significantly changed.
Appendix G Prompt
G.1 Prompt Template for Query Grounding
Here is the prompt we used for query grounding.
Prompt Template for Query Grounding Analyze the following video frames and the question: Question: <Question> Options: <Options> Step 1: Key Object Identification • Extract 3-5 core objects detectable by computer vision • Use YOLO-compatible noun phrases (e.g., “person”, “mic”) • Format: Key Objects: obj1, obj2, obj3 Step 2: Contextual Cues • List 2-4 scene elements that help locate key objects based on options provided • Use detectable items (avoid abstract concepts) • Format: Cue Objects: cue1, cue2, cue3 Step 3: Relationship Triplets • Relationship types: • Spatial: Objects must appear in the same frame • Attribute: Color/size/material descriptions (e.g., “red clothes”, “large”) • Time: Appear in different frames within a few seconds • Causal: There is a temporal order between the objects • Format of Relations: (object, relation_type, object), relation_type should be exactly one of spatial/attribute/time/causal Output Rules 1. One line each for Key Objects/Cue Objects/Rel starting with exact prefixes 2. Separate items with comma except for triplets where items are separated by semicolon 3. Never use markdown or natural language explanations 4. If you cannot identify any key objects or cue objects from the video provided, please just identify the possible key or cue objects from the question and options provided Below is an example of the procedure: Question: For “When does the person in red clothes appear with the dog?” Response: Key Objects: person, dog, red clothes Cue Objects: grassy_area, leash, fence Rel: (person; attribute; red clothes), (person; spatial; dog) Format your response EXACTLY like this in three lines: Key Objects: object1, object2, object Cue Objects: object1, object2, object Rel: (object1; relation_type1; object2), (object3; relation_type2; object4)
G.2 Prompt Template for Question Answering
Here is the prompt we used for question answering.
Prompt Template for Question Answering Select the best answer to the following multiple-choice question based on the video. <image> <image> $\cdots$ Question: <Question> Options: <Options> Answer with the option’s letter from the given choices directly. Your response format should be strictly an upper case letter A,B,C,D or E.