### v2

# Logic-in-Frames: Dynamic Keyframe Search via Visual Semantic-Logical Verification for Long Video Understanding

Abstract
Understanding long video content is a complex endeavor that often relies on densely sampled frame captions or end-to-end feature selectors, yet these techniques commonly overlook the logical relationships between textual queries and visual elements. In practice, computational constraints necessitate coarse frame subsampling, a challenge analogous to “finding a needle in a haystack.” To address this issue, we introduce a semantics-driven search framework that reformulates keyframe selection under the paradigm of Visual Semantic-Logical Search. Specifically, we systematically define four fundamental logical dependencies: 1) spatial co-occurrence, 2) temporal proximity, 3) attribute dependency, and 4) causal order. These relations dynamically update frame sampling distributions through an iterative refinement process, enabling context-aware identification of semantically critical frames tailored to specific query requirements. Our method establishes new SOTA performance on the manually annotated benchmark in key-frame selection metrics. Furthermore, when applied to downstream video question-answering tasks, the proposed approach demonstrates the best performance gains over existing methods on LongVideoBench and Video-MME, validating its effectiveness in bridging the logical gap between textual queries and visual-temporal reasoning. The code will be publicly available.
1 Introduction
Vision-Language Models (VLMs) yin2024survey have achieved remarkable progress in video understanding zou2024seconds; tang2023video, particularly in video question answering wang2024gpt4video; 84, demonstrating potential for modeling real-world scenarios. However, existing methods can only simultaneously process a limited number of frames due to the inherent token limit and extremely high dimension of spatio-temporal video data, especially for long videos. Furthermore, uniformly sampled keyframes are query-agnostic and insufficient to represent query-related contents. To tackle these challenges, this paper addresses a pivotal research question:
How can we efficiently and accurately select keyframes that are semantically critical for answering video-based queries?
We hypothesize that deconstructing visual semantic and logical cues (e.g., target objects, logical relations including temporal, spatial, attribute, and causal relationships between visual entities) from textual queries enables effective identification of task-relevant frames through heuristic sampling and search. Building on this insight, we propose Visual Semantic-Logical Search (VSLS), a novel keyframe search framework that incorporates target object confidence estimation and joint verification of visual semantic logic into the iterative update of frame sampling distribution and selects the most informative frames with the highest confidence. Experimental results show that our approach requires only sparse sampling (1.4% of frames per video on average) to identify critical frames, significantly reducing computational complexity compared to conventional dense sampling strategies while maintaining performance on downstream video understanding tasks.
Figure 1: Examples of four types of visual semantic-logical relationships in video QA detected by our VSLS framework: Temporal (text, time, pen), Attribute (man, attribute, white shirt), Spatial (copilot, spatial, Egyptian Pyramids), and Causal (man, causal, basketball). Green boxes indicate correct answers, while red boxes show baseline errors.
Compared to conventional methods, VSLS shows three distinct advantages. First, the framework is training-free and highly efficient in comparison with dense captioning chen2024sharegpt4video; kim2024you; wang2024videoagent or video clustering wang2024videotree; rajan2025key strategies, sampling only 1.4% of frames on average in LVHaystack. Second, it explicitly models logical binary relations (namely spatial, temporal, attribute, and causal) in the query beyond simple target detection ye2025rethinking, utilizing additional visual semantic features and enhancing logical consistency throughout the reasoning process. Third, VSLS is a plug-and-play module, which can be seamlessly integrated into existing VLM pipelines without cross-component dependencies.
We further examine VSLS on several public datasets, including LongVideoBench lvhdataset, a comprehensive benchmark for long video understanding; Video-MME videomme, a widely adopted multimodal video question answering dataset; and Haystack-LVBench lvhdataset with meticulously annotated keyframes based on human feedback for more precise analysis. Extensive experiments demonstrate significant improvements in both the semantic similarity and temporal coverage between the retrieved keyframes and the ground truth labels, as well as the accuracy in downstream video question-answering tasks. More importantly, with only 1.4% of video frames (Ego4D grauman2022ego4d) sampled in the search iteration, our method achieves an 8.7% improvement in GPT-4o hurst2024gpt’s long video QA accuracy. This performance gain is attributed to our simple yet powerful observation: query-guided visual semantic logic retrieval can mitigate the gap between potential visual logic in video frames and the logic expressed in the query. To be specific, constructing ternary logic triplets with visual elements (e.g., object1, logic type, object2) can enhance downstream reasoning capabilities when performing textual-visual retrieval.
To the best of our knowledge, we are arguably the first to search for keyframes in long videos by detecting visual semantic logic, with potential extensions to other textual-visual retrieval tasks. Our main contributions are as follows:
  • We define four fundamental types of semantic logic relations in video QA tasks, including temporal, causal, attribute, and spatial relations, which can be accurately detected across various datasets.
  • We sample only 1.4% of frames on average of frames on average during keyframe search through heuristic sampling and distribution updating by different visual semantics and logical relations.
  • We comprehensively evaluate retrieval efficiency, semantic similarity, temporal coverage, and video question answering accuracy across several widely used video understanding datasets, demonstrating significant improvements in downstream tasks.
2 Method
Figure 2: Our VSLS Framework for Efficient Keyframe Selection. VSLS sparsely samples frames and selects key ones via object detection and logic verification. Steps: 1) Use LLM&VLM to extract cue/target objects and four logic types (spatial, temporal, attribute, causal); 2) Adaptive sampling with evolving confidence; 3) Detect objects via YOLO-World; 4) Fuse scores with a spline function to identify high-confidence frames for downstream tasks.
Although existing long-context VLM frameworks implement keyframe search for video QA tasks 35; 47; 57; 65; 68; 80, their computational efficiency and searching accuracy remain suboptimal. To address this needle-in-a-haystack challenge 61; zhao2024needle, we propose a novel method VSLS that aligns the semantic relations between the text modality and video modality, enhancing the plausibility of logical reasoning and performance of downstream tasks.
2.1 Task Formulation
Given a video sequence $V=\{f_{t}\}_{t=1}^{N_{v}}$ with $N_{v}$ frames and a query $Q$, the ideal temporal search framework aims to retrieve the minimal keyframe subset $V^{K}=\{f_{m_{i}}\}_{i=1}^{K}\subseteq V$ with $K$ keyframes that satisfies:
  • Conservation: The keyframe subset $V^{K}\subseteq V$ must satisfy the answer consistency condition: $\mathcal{A}(V^{K},Q)=\mathcal{A}(V,Q)$, where $\mathcal{A}(\cdot)$ denotes the video QA function.
  • Compactness: $V^{K}$ must be a minimal subset that preserves completeness, which means that no frame in $V^{K}$ can be removed without hindering the accuracy and efficiency of video QA.
2.2 Visual Semantic Logic Extraction
Starting from a question $Q$ and uniformly sampled frames $\overline{V}_{N}$ from video $V$, our goal is to extract key visual elements to answer $Q$. We first classify the detected objects in $Q$ and $\overline{V}_{N}$ into two categories:
  • Key Objects: The main participants or references in the scene that the question explicitly or implicitly focuses on (e.g., “person”, “microphone”).
  • Cue Objects: Secondary or contextual entities that help locate or disambiguate the Key Objects (e.g., “book”, “tiger painting”).
To further leverage semantic and logical links among these objects, we define a set of relations $\mathcal{R}\subseteq\mathcal{O}\times\Delta\times\mathcal{O}$, where each relation $r=(o_{i},\delta,o_{j})\in\mathcal{R}$, with $o_{i},o_{j}\in\mathcal{O}$ denoting detected objects in the key and cue objects dataset, and $\delta\in\Delta$ representing one of the following types of relations:
[…]
$o_{i}$ and $o_{j}$ appear in the same frame, indicating co-occurrence or proximity.
Example: “A person is standing beside a vase.” $\Rightarrow$ (person, spatial, vase)
$o_{i}$ and $o_{j}$ share visual properties, e.g., color or size.
Example: “A person wears a black shirt.” $\Rightarrow$ (person, attribute, black shirt)
[…]
$o_{i}$ and $o_{j}$ occur in close frames, linking sequences or transitions.
Example: “After a dog entered the room, a cat entered.” $\Rightarrow$ (dog, temporal, cat)
$o_{i}$ and $o_{j}$ follow a cause-effect or prerequisite order.
Example: “A little girl broke the vase.” $\Rightarrow$ (little girl, causal, pieces)
The choice of these four relations draws on core concepts in linguistics and logic cohen1968universals; sowa2000knowledge; talmy2000toward, which identify spatial, temporal, attributive, and causal aspects as fundamental for structuring, perceiving, and communicating information about events and states. For more details on this selection, please see appendix LABEL:Theoretical for reference. As shown in Figure 1, we construct semantic-logical relations that support a broad range of question-answering tasks. Specifically, questions involving temporal queries (when does X happen?”), causal reasoning (why did Y occur?”), attribute dependence (What is the person wearing sunglasses doing?”), or spatial constraints (Who is standing next to the red car?”) can be answered more reliably by incorporating these structured relations and contextual cues.
2.3 Iterative Semantic-Logical Temporal Search
Based on the extracted key and cue objects and their logic relations, our algorithm iteratively searches for keyframes through semantic and logical reasoning, including four main stages: Frame Sampling (Sec. 2.3.1), Object Detection and Scoring (Sec. 2.3.2), Visual Semantic Logic Detection (Sec. 2.3.3), and Distribution Update (Sec. 2.3.4). The pseudocode is shown in Algorithm 1, and Algorithm LABEL:alg:slt_search provides a more detailed version.
Function SemanticLogicalTemporalSearch($V,Q,K,\Delta_{t},\tau,\alpha,\gamma$)
       $\mathcal{O},\mathcal{R}\leftarrow\text{ParseQuestion}(Q)$ // Extract key/cue objects and relations
      $P\leftarrow\text{Uniform},B\leftarrow|V|,S\leftarrow\emptyset,N_{v}\leftarrow|V|$ // Initialize distribution and state
      while $B>0$ and $|\mathcal{O}|>0$ do
             $k\leftarrow\lfloor\sqrt{B}\rfloor$, $G\leftarrow\textnormal{{Grid}}(\textnormal{{Sample}}(P,k^{2}))$ // Adaptive grid sampling
            $\Omega\leftarrow\textnormal{{DetectObjects}}(G)$ // Detect objects in sampled frames
            foreach $t\in G$ do
                   $C_{t}\leftarrow\textnormal{{CalculateBaseScore}}(\Omega_{t})$ // Base detection confidence
                  foreach  $r_{type}\in\mathcal{R}$  do
                         $\delta\leftarrow\textnormal{{Processrelation}}(r_{type},\Omega,\Delta_{t},\tau,\alpha,\gamma)$ // relations require distinct processing
                        $C_{t}\leftarrow C_{t}+\delta$
                  $\textnormal{{UpdateScores}}(S,t,C_{t})$ // Update global score registry
            $\textnormal{{DiffuseScores}}(S,w)$ // Temporal context propagation
            $P\leftarrow\textnormal{{NormalizeDistribution}}(S)$, $B\leftarrow B-k^{2}$ // Update sampling distribution
            foreach $g\in\textnormal{{TopK}}(S,K)$ do
                   if $\Omega[g]\cap\mathcal{O}\neq\emptyset$ then
                         $\mathcal{O}\leftarrow\mathcal{O}\setminus\Omega[g]$ // Remove identified key objects
      return $\textnormal{{TopK}}(S,K)$ // Return top-K keyframes
Algorithm 1 Visual Semantic-Logical Search
2.3.1 Frame Sampling
To accelerate the search process, we avoid exhaustively scanning all $N_{v}$ video frames and instead employ a distributed sampling strategy. Let $N_{v}$ denote the total number of frames in the video, and $P$ be a uniformly initialized sampling distribution over all frames. The sampling process is then defined as:
$I_{s}=\mathrm{Sample}(P\odot N_{v},N_{s}),$ (1)
where $\mathrm{Sample}(\cdot,N_{s})$ selects a subset of $N_{s}$ frames according to the distribution $P\odot N_{v}$. To further leverage the detecting ability of YOLO, we stack the sampled frames into a $k\times k$ grid, which imposes a constraint on the sample size $N_{s}$. Specifically, we require:
$N_{s}\in\{k^{2}\mid k\in\mathbb{Z}\}\quad\text{and}\quad N_{s}<N_{v}.$ (2)
In practice, this ensures that the number of sampled frames can be reshaped into a compact 2D grid for efficient processing. Although $P$ is initially uniform, it can be adapted over multiple rounds of sampling to focus on frames of higher interest in the video.
2.3.2 Object Detection and Scoring
In this stage, we construct the detection search space by taking the union of both key objects and cue objects. For each iteration, we detect objects on the $N_{s}$ sampled frames using a lightweight model like YOLO-World yolo for high efficiency and score the frames based on detection confidence. Specifically, let $\Omega_{t}$ be the set of detected objects in the frame at time $t$, $c_{o}$ the confidence of each detected object, and $w_{o}$ the corresponding weight. We define the frame score as:
$C_{t}=\max_{o\in\Omega_{t}}\bigl{(}c_{o}\cdot w_{o}\bigr{)}.$ (3)
If the confidence score of any key object exceeds a predefined threshold, it is added to a list, thereby maintaining a record of frames where crucial targets have been identified for subsequent processing.
2.3.3 Visual Semantic Logic Detection
Figure 3: Sample weight evolution under VSLS optimization for keyframe selection. Top: 16 iterations show progressive convergence toward Ground Truth (red). Bottom: 15 iterations demonstrate similar alignment. Yellow highlights indicate precise matches between algorithm outputs (green) and manual annotations.
Beyond individual object detection and frame-level scoring, we refine each frame’s confidence score by modeling higher-order object relations. Let $\mathcal{R}$ be the set of relations, where each $r\in\mathcal{R}$ involves a pair $(o_{1},o_{2})$ and is labeled by a type $r_{\text{type}}$. Denote $C_{t}$ as the confidence score at time $t$, with a global scaling factor $\alpha$ and a relation-specific weight $\gamma_{r_{\text{type}}}$ controlling each logic type’s impact. The refined confidence $C_{t}^{(r)}$ after applying relation $r$ is defined as:
$C_{t}^{(r)}=C_{t}+\alpha\cdot\gamma_{r_{\text{type}}}.$ (4)
Spatial Relation. A spatial relation enforces that two objects $o_{1}$ and $o_{2}$ must co-occur in the same frame. Let $\Omega_{t}$ be the set of detected objects in frame $t$. If both $o_{1}\in\Omega_{t}$ and $o_{2}\in\Omega_{t}$, then the corresponding frame confidence is updated as:
$C_{t}\leftarrow C_{t}+\alpha\cdot\gamma_{\text{spatial}}.$ (5)
Attribute Relation. An attribute relation is satisfied when $o_{1}$ and $o_{2}$ share sufficient bounding-box overlap in the same frame. Let $\mathrm{overlap}$ be the ratio of their intersection area to the minimum of their individual bounding-box areas. If the overlap ratio exceeds a predefined threshold $\tau$ ($\tau=0.5$ in our experimental setting), we increase the frame confidence:
$C_{t}\leftarrow C_{t}+\alpha\cdot\gamma_{\text{attribute}}.$ (6)
Time Relation. A time relation checks whether two objects appear in temporally close frames. Suppose $t_{i}$ and $t_{j}$ ($t_{i}\leq t_{j}$) are sampled such that $|\,t_{j}-t_{i}|<\Delta_{t}$, where $\Delta_{t}$ is a threshold (e.g. 5 frames in our experimental setting), if $o_{1}$ occurs in frame $t_{i}$ and $o_{2}$ in frame $t_{j}$, then both frames’ confidences are updated:
$C_{t_{i}}\leftarrow C_{t_{i}}+\alpha\cdot\gamma_{\text{time}},\quad C_{t_{j}}\leftarrow C_{t_{j}}+\alpha\cdot\gamma_{\text{time}}.$ (7)
Causal Relation. A causal relation models an ordering constraint, enforcing that $o_{1}$ must appear at an earlier time than $o_{2}$. Specifically, if $o_{1}\in\Omega_{t_{i}}$ and $o_{2}\in\Omega_{t_{j}}$ with $t_{i}<t_{j}$, we update the confidences of frames $t_{i}$ and $t_{j}$ by:
$C_{t_{i}}\leftarrow C_{t_{i}}+\alpha\cdot\gamma_{\text{causal}},\quad C_{t_{j}}\leftarrow C_{t_{j}}+\alpha\cdot\gamma_{\text{causal}}.$ (8)
Through this scoring mechanism, frames with detected relations will have greater confidence and are more likely to be retrieved as keyframes for the given query and video. We have also conducted hyperparameter search experiments, and find that $\alpha=0.3$ (from 0.3, 0.5, 0.7, 1.0) and $\gamma_{r_{\text{type}}}=0.5$ achieve the best results across different datasets.
2.3.4 Distribution Update
After each iteration of frame sampling, we merge newly obtained frame confidences into the global score distribution $\{S_{f}\}$ spanning all frames $f=1,2,\ldots,N_{v}$. When a frame $f$ is selected for detection, its score is assigned to the confidence value $C_{f}$, and the visitation counter $N_{v,f}$ is reset to 0. To incorporate temporal context, we diffuse this updated score to neighboring frames within a window of size $w$. Denoting each nearby index by $f\pm\delta$ (for $\delta\in[-w,w]$), we apply:
$S_{f\pm\delta}\leftarrow\max\left(S_{f\pm\delta},\,\frac{S_{f}}{1+|\delta|}\right).$ (9)
In this way, high-confidence frames raise the scores of close-by frames, reflecting temporal continuity. Following these local updates, the sampling distribution $P$ is refined using spline interpolation, and then normalized. This iteration proceeds until either the search budget $B$ is reached or all key objects have been successfully identified. The visualization of distribution in different iterations can be seen in Figure 3. Finally, the method outputs the top $K$ frames according to their terminal scores.
3 Experiment
3.1 Benchmark Datasets
The proposed VSLS is systematically evaluated across four benchmark datasets: a) LongVideoBench lvhdataset for assessing long-context video-language comprehension capabilities; b) Video-MME videomme as the first comprehensive benchmark for multimodal video analytics; c) Haystack-LVBench, extended from LongVideoBench with human-annotated frame index answers; and d) Haystack-Ego4D, derived from Ego4D with similar annotations. While LongVideoBench and Video-MME measure performance enhancement in QA accuracy, Haystack-Ego4D and Haystack-LVBench quantitatively evaluate keyframe selection accuracy through recall and precision metrics. Further details of datasets are provided in Appendix LABEL:app:details.
3.2 Evaluation Metrics
3.2.1 Evaluation Metrics for Search Utility
Our assessment framework emphasizes both effectiveness and efficiency. For search effectiveness, we use three metrics to compare model-predicted keyframes with human annotations, considering both individual frames and full sets—addressing the possibility of multiple valid keyframe sets per query. For frame-level comparison, we evaluate the alignment between a predicted frame $f_{\text{pt}}$ and a human-annotated frame $f_{\text{gt}}$ from two perspectives:
Temporal coverage evaluates the coverage of ground truth frames by predicted frames in the temporal perspective, which can be described as:
$T_{\text{cover}}(T_{\text{pt}},T_{\text{gt}})=\frac{\sum\limits_{i=1}^{|N_{\text{gt}}|}\mathbb{I}\left[\min\limits_{j}\left|t_{\text{gt}}^{i}-t_{\text{pt}}^{j}\right|\leq\delta\right]}{|N_{\text{gt}}|},$ (10)
where $T_{\text{pt}}$ and $T_{\text{gt}}$ denote the sets of predicted and ground truth timestamps, respectively. Here, $|N_{\text{gt}}|$ is the number of ground truth frames, $t_{\text{gt}}^{i}$ and $t_{\text{pt}}^{j}$ are the $i$-th ground truth and $j$-th predicted timestamps, respectively. $\delta$ is the temporal similarity threshold defining the maximum allowed time deviation, and $\mathbb{I}[\cdot]$ is the indicator function, returning 1 if the condition holds and 0 otherwise.
Visual Similarity is measured by the Structural Similarity Index (SSIM) ssim, capturing structural detail, luminance, and contrast between $f_{\text{pt}}$ and $f_{\text{gt}}$. For set-to-set comparison, the key challenge is defining inter-set similarity. We adopt Precision $P$ and Recall $R$ as complementary metrics: Precision checks whether each predicted frame matches any reference frame, while Recall ensures that all reference frames are represented. Given the ground truth set $F_{\text{gt}}={f^{j}{\text{gt}}}^{n}{j=1}$ and the predicted set $F_{\text{pt}}={f^{i}{\text{pt}}}^{m}{i=1}$, we define the multimodal retrieval quality metrics as follows:
$P(F_{\text{pt}},F_{\text{gt}})$ $=\frac{1}{|F_{\text{pt}}|}\sum_{f^{i}_{\text{pt}}\in F_{\text{pt}}}\max_{f^{j}_{\text{gt}}\in F_{\text{gt}}}\phi(f^{i}_{\text{pt}},f^{j}_{\text{gt}}),$ (11a)
$R(F_{\text{pt}},F_{\text{gt}})$ $=\frac{1}{|F_{\text{gt}}|}\sum_{f^{j}_{\text{gt}}\in F_{\text{gt}}}\max_{f^{i}_{\text{pt}}\in F_{\text{pt}}}\phi(f^{j}_{\text{gt}},f^{i}_{\text{pt}}),$ (11b)
where $\phi(\cdot,\cdot)$ represents an extensible multimodal similarity metric function.
3.2.2 Evaluation Metrics for Search efficiency
Existing studies 13; 47; 65; 68; 71 have mainly concentrated on optimizing task-specific performance metrics while neglecting computational efficiency in temporal search operations. To systematically analyze this dimension, our evaluation framework incorporates two criteria: 1) FLOPs representing arithmetic operation complexity, and 2) Latency recording real-world execution duration.
3.3 Evaluation of Search Framework efficiency
[…]
[gray]0.7
[…]
Table 1: Evaluation of performance metrics across the LV-Haystack benchmark, presenting both search efficiency and end-to-end processing overhead (combining search and inference stages).
Current approaches for keyframe selection can be broadly categorized into three paradigms: statistic-based frame sampling, dense feature retrieval-based selection, and temporal search-based methods. As shown in Table 1, while uniform sampling achieves the fastest processing speed, its ignorance of frame semantics severely limits downstream task effectiveness. Although dense feature retrieval methods attain moderate accuracy improvements (57.3%), their exhaustive frame processing demands 4.2$\times$ more TFLOPs and introduces 4.5$\times$ higher latency than our temporal search approach. Crucially, our method introduces four visual semantic logic detectors during temporal search while maintaining comparable execution time to T* methods. This strategic design elevates downstream task accuracy to 61.5%, achieving the best performance-efficiency trade-off.
3.4 Visual Semantic Logic Search Performance
As demonstrated in Table 2, we evaluate VSLS on LongVideoBench from two critical perspectives: visual similarity (measured by precision and recall) and temporal coverage. Our method achieves state-of-the-art performance across all metrics. Specifically, under the 32-frame setting, VSLS attains a precision of 74.5% and recall of 92.5%, outperforming all baselines in visual similarity. More notably, the temporal coverage of VSLS reaches 41.4%, surpassing the second-best method (T* at 36.5%) by 13.4%—the largest margin among all comparisons. This significant improvement highlights the effectiveness of our visual semantic logic detection modules in identifying query-relevant keyframes with both semantic alignment and temporal completeness.
Table 2: Search utility results on LongVideoBench. Best scores in the 8-frame setting are underlined, and in the 32-frame setting are bold. Gray indicates results from the original paper.
[…]
These results empirically support our core hypothesis: leveraging semantic and logical cues from text queries enables precise detection of relevant video frames. Improvements in visual similarity and temporal coverage confirm that VSLS effectively captures keyframes while preserving temporal coherence through visual-logical alignment.
3.5 Downstream Video QA Performance
[…]
[gray]0.7